# A general approach to modeling environmental mixtures with multivariate outcomes

### v1

@@ 5.1 Simulation A: Distributed lag models with multiple exposures
In Figure 2 we report MSE for $\bm{\omega}_{kj}$ when $n=500$. For brevity, we highlight results for the first two outcomes and first three exposures (with other results being similar). Clustering generally improved estimation of $\bm{\omega}_{kj}$: models with more clustering (DLNM, DLNM dimension reduction) had lower MSE than others. The benefits were largest for exposure 1 as it shared lag profiles with exposures 4 and 5, hence allowed for more borrowing of information. We also observe a tradeoff: whereas the DLNM with dimension reduction dramatically improved computation time, it yielded slightly higher MSE than without. We found similar results for estimation of $f_{kj}(\cdot)$, with some minor differences (see Appendix D). While clustering was generally beneficial for estimation of $f_{kj}(\cdot)$, clustering only across exposures (DLNM separate) outperformed full clustering for outcomes 1 and 3. As mentioned above, this is because $f_{j1}(\cdot)$ and $f_{j^{\prime}3}$ were both linear but with different slopes, and the DLNM occasionally incorrectly clustered them together—but again this improved substantially with higher sample size (Appendix D.3).
Figure 2: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators. MSE values are scaled so that DLNM has an MSE of 1 in all scenarios.
Figure D9 in Appendix D shows that all four models yielded near nominal 90% coverage for 16 of the 20 exposure-outcome pairs, though coverage was somewhat lower for $(k,j)=$ (1,2) and (1,3) (and analogously for outcome 3). Lag weights $\bm{\omega}_{kj}$ for these pairs were commonly estimated to be flat as a result of being incorrectly clustered with other exposures for outcome 1. This shows the bias-variance trade-off inherent to clustering: MSE was generally lower for the clustering models, but this came from a significant reduction in variance and a slight increase in bias (which can affect coverage). As mentioned above, clustering improved in larger samples, and this led to better coverage. We note that this problem of under-coverage for these lag weights occurred even in the absence of clustering (albeit less so) due to the weak identifiability of parameters.
@@ 5.2 Simulation B: Single time point mixture modeling
Now we shift focus to settings where the exposures are only observed at a single time point, but index models can be used to model non-linear effects of exposures on the outcomes.
We generated datasets with $K$=4 outcomes and $P$=10 exposures and varied the sample size $n\in\{200,600,1000\}$. The exposures $\bm{x}$ were generated from a multivariate normal distribution with a moderate autoregressive correlation structure ($\rho$=0.5). Exposure-outcome associations were weak to moderate, with signal to noise ratios between 0.07 and 0.37, and we considered three scenarios that vary in their favorability to MIMs:
Scenario 1: $E[\bm{Y}]=\left[\begin{array}[]{c}0.5f_{1}(\bm{x})\\ 0.5f_{1}(\bm{x})\\ 0.5f_{2}(\bm{x})\\ 0.3f_{2}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=\bm{x}^{T}\bm{\alpha}_{1},$
$f_{2}(\bm{x})=\bm{x}^{T}\bm{\alpha}_{1}+(\bm{x}^{T}\bm{\alpha}_{2})^{2},$
@@
Scenario 2: $E[\bm{Y}]=\left[\begin{array}[]{c}0.25f_{1}(\bm{x})\\ 0.25f_{1}(\bm{x})\\ 0.3f_{2}(\bm{x})\\ 0.3f_{2}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=e^{0.5x_{1}}+\dfrac{1.5e^{1.2x_{2}}}{1+e^{1.2x_{2}}}-0.5x_{3}^{2},$
$f_{2}(\bm{x})=x_{9}-0.75x_{10}^{2}.$
Scenario 3: $E[\bm{Y}]=\left[\begin{array}[]{c}f_{1}(\bm{x})\\ f_{1}(\bm{x})\\ \sqrt{0.5}f_{1}(\bm{x})\\ \sqrt{0.5}f_{1}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=(\bm{x}^{T}\bm{\alpha}_{1})\times(\bm{x}^{T}\bm{\alpha}_{2}),$
@@
To each dataset we fit four different mixtures approaches:
  1. 1.
    The proposed MIM applied separately to each outcome (MIM separate)
  2. 2.
    The proposed multivariate MIM, which clusters across outcomes (MIM clustering)
  3. 3.
    Bayesian kernel machine regression (BKMR, Bobb et al.,, 2015), a widely used approach in environmental epidemiology that models potentially non-linear and non-additive relationships using a Gaussian process that also performs variable selection on the exposures.
  4. 4.
    Weighted quantile sum regression (WQS, Carrico et al.,, 2015), a popular mixtures approach that fits a linear index model on exposure quantiles.
@@
We also computed exposure importance metrics $\Psi_{kj}$ proposed in Section 3.1 for the MIM. To estimate the distribution of $x_{j}$ given $\bm{x}_{-j}$ we assumed normality and estimated the mean of $x_{j}$ given $\bm{x}_{-j}$ via random forests to allow for flexible associations between exposures. We report MSE for the posterior mean of the exposure-response surface averaged over 200 randomly chosen locations, and we also report the average exposure importance metrics.
@@
We report results for MSE in Figure 3. In Scenario 1, the true exposure-response relationships fit into the index framework, and the MIM approaches naturally performed best. By contrast, Scenario 2 is favorable to BKMR, as the regression surfaces are sparse (70-80% of effects are null), and each non-null effect has a different functional form. As such, BKMR performed best here in small sample sizes. Surprisingly, however, the proposed MIM approaches began to outperform BKMR in larger samples, as BMKR tended to incorrectly exclude important exposures in some datasets. Scenario 3, characterized by many pairwise interactions, is not particularly favorable to either approach. Here the MIMs—despite having an additive structure—greatly outperformed the other approaches. Across all scenarios and sample sizes, MIM clustering successfully exploited shared structure across outcomes, and outperformed MIM separate (which did not borrow information across outcomes). Overall, the proposed approach fared well in a range of settings: under sparsity, it performed reasonably well, and when many exposures had small effects, it greatly outperformed competitors.
@@
Figure 3: Mean squared error for estimating $f^{*}(\bm{x})$ at 200 randomly chosen locations for $\bm{x}$ across all estimators, sample sizes, and scenarios. MSE values are scaled so that MIM clustering has an MSE of 1 in all scenarios.
@@
In Figure 4, we see that estimation of exposure importance metrics is difficult in small sample sizes. In particular when $n=200$ we observe substantial overestimation in the highly sparse Scenario 2, in part because when $\Psi_{kj}=0$ it can never be underestimated (as $\Psi_{kj}\geq 0$ by definition). Ultimately, as sample size increased, estimation improved significantly in all three scenarios. This shows that the proposed exposure metrics can be reasonably estimated in order to provide additional interpretability to the proposed MIM framework.
@@ 6 NMMAPS Data Analysis
The National Morbidity, Mortality and Air Pollution Study (NMMAPS) collected time series data on daily mortality, air pollution and meteorological variables in Chicago between 1987 and 2000. We observe daily cardiovascular deaths, respiratory deaths and other non-accidental deaths, as well as median PM10 and O3 measurements. Full study details can be found in Samet et al., 2000a ; Samet et al., 2000b . Mortality outcomes were log transformed and standardized. Exposures were standardized, and we excluded one two week period in 1988 with a PM10 measurement 15 standard deviations above the mean, leaving a dataset of $n=5,100$ days.
@@
To estimate the associations between exposure and the three outcomes, we fit the proposed MV-DLNM with $K=3$ and $P=2$, with a maximum time-lag of $L=14$ days. Following Chen et al., (2019), we adjusted for day of week (categorical), seasonality (via cyclic basis splines for month, dim=4), year (via thin plate splines, dim=6), and mean daily temperature and dew point temperature both per day (via thin plate splines, dim=6) as well as averaged over the previous three days (via thin plate splines, dim=3). We allowed for different covariate associations for each outcome ($\bm{\beta}_{Z1}\neq\bm{\beta}_{Z2}\neq\bm{\beta}_{Z3}$). We fit the MV-DLNM with and without clustering to investigate the effect of borrowing strength across outcomes and exposures.
Directly comparing exposure-response curves $f_{kj}(\cdot)$ across different models can be non-intuitive, since they are functions of different estimates of $\bm{\omega}_{kj}$. Instead, similar to Chen et al., (2019), we report mean outcome differences corresponding to an increase in a pollutant from 1SD below its mean to 1SD above at each lag $l$, holding exposures at all other lags constant. We highlight results for PM10; associations for ozone are all close to null with high uncertainty (see Appendix E). Results for PM10 without clustering are reported in the left hand column of Figure 5. Associations were small across all three outcomes, because these contrasts correspond to changes in a single day exposure while holding the other 13 days’ exposures constant. Estimates varied somewhat across different outcomes. For respiratory deaths, the strongest associations occurred at lag-0, and these decayed smoothly to the null after two weeks. Associations were also strongest at lag-0 for other non-accidental deaths, although by the end of the lag window associations had the opposite sign. By contrast, the association with cardiovascular deaths was null at lag-0 but peaked at lags 7–9.
@@
Figure 5: Mean difference in outcomes increasing PM10 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is without clustering; right column is with clustering.
Estimating $K\times P=6$ non-parametric functions $f_{kj}(\cdot)$, which are themselves functions of $6$ different unknown weight vectors $\bm{\omega}_{kj}$ of length $L=14$ makes identifying small effects challenging, so we hypothesized that the proposed clustering strategy would improve efficiency. In Figure 6 we report pairwise posterior clustering probabilities for $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$. The degree of clustering for $f_{kj}(\cdot)$ was extremely high, indicating similar associations for all outcome-exposure pairs except for the cardiovascular - O3 relationship. There was less clustering in the time-lag profiles, but results still indicate substantial similarity, especially for PM10. This is borne out by estimated associations in Figure 5 (right hand column), in which estimates are quite similar across different outcomes, with the strongest effects at lag-0 and decaying to null by the end of the lag window. In addition to improved stability, it also yielded narrower CIs, especially for respiratory deaths. Also note that the counter-intuitive negative association at time-lag of 13 in the model is removed once clustering is incorporated.
Figure 6: Pairwise posterior clustering probabilities for Chicago analysis. Left plot is for $\bm{\beta}$; right plot for $\bm{\theta}$
@@ 7 Discussion
In this paper we have proposed a general framework for modeling environmental mixtures and possibly multivariate outcomes. Although we highlighted the distributed lag framework for time-varying exposures, the proposed approach readily applies to other exposure structures. Rather than measuring $P$ pollutants at $L$ different times, for example, Bauer et al., (2020) analyzed five different biomarkers (in hair, blood, urine, nails, and saliva) of four different metals, for each of three different outcomes via separate BKMR models. Our approach can be applied to this setting by modeling each of the $P\times K=4\times 3$ outcome-exposure relationships via $f_{kj}(\bm{x}_{ij}^{T}\bm{\omega}_{kj})$, where $\bm{x}_{ij}$ is a vector of length $L=5$ representing the $p^{th}$ metal measured via each of the five different biomarkers, and $\bm{\omega}_{kj}$, representing the relative importance of each of these biomarker measurements to the overall association. Even with a mixture of univariate exposure measurements, the proposed framework extends additive models by encouraging similar exposure-response functions across different exposures and outcomes:
$y_{ik}$ $=f_{k1}({{x}}_{i1})+\cdots+f_{kj}({{x}}_{iP})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~{}~{}k=1,\dots,K.$
Here the weights are fixed ($\omega_{kj}=1$), but we can still improve performance by encouraging similarity in the functions $f_{kj}(\cdot)$. Naturally, we expect the greatest gains when $P$$\times$$K$ is large, but the proposed approach may still offer a benefit even if the outcome is univariate ($K=1$) by borrowing strength across a large number $P>1$ of exposures. Similarly, with a univariate exposure ($P=1$), the proposed approach can borrow strength across $K>1$ outcomes.
Index models offer another approach to improving power in mixtures analyses by imposing more structure than fully non-parametric approaches like BKMR. In contrast to more flexible approaches, MIMs model the exposure-response surface as a function of $M<P$ linear indices, thus reducing the dimensionality of the estimation task. When the true number of indices—or their correct compositions—is unknown, however, they can yield bias (McGee et al.,, 2021). As an extension of our development, we have shown how to fit (possibly multivariate) MIMs that are robust to misspecification of the true index structure.
Software and Data Availability
Data analyzed in Section 6 are available in the R package dlnm Gasparrini, (2011). R code for running the proposed methods are also available at github.com/glenmcgee/MVmixture.

### v2

@@ 5.1 Simulation A: Distributed lag models with multiple exposures
In Fig. 2 we report MSE for $\bm{\omega}_{kp}$ when $n=500$. We highlight results for the first two outcomes and first three exposures (with others being similar). Clustering generally improved estimation of $\bm{\omega}_{kp}$: models with more clustering (DLNM, DLNM dimension reduction) had substantially lower MSE than others. We also observe that the DLNM with dimension reduction not only improves computation time, but can at times improve estimation error as well, likely due to the reduced number of parameters being estimated. We found similar results for estimation of $f_{kp}(\cdot)$, where clustering generally provides much lower mean squared error (see Appendix G). Appendix Fig. G9 shows that all four models yielded near nominal coverage for most of the exposure-outcome pairs, though coverage was lower for $(k,p)=$ (2,2) and (3,2). Lag weights $\bm{\omega}_{kp}$ for these pairs were commonly estimated to be flat as a result of being incorrectly clustered with other exposure-outcome pairs. This shows the bias-variance trade-off inherent to clustering: MSE was generally lower for the clustering models, but this came from a significant reduction in variance and a slight increase in bias. In larger samples, clustering improved and led to nominal coverage rates.
Figure 2: Mean squared error for estimating $\bm{\omega}_{kp}$ across all estimators. MSE values are scaled so that DLNM has an MSE of 1 in all scenarios.
@@ 5.2 Simulation B: Single time point mixture modeling
We now consider a setting where exposures are observed at a single time point as in Section 4. We generated datasets with KK=4 outcomes and PP=10 exposures and varied the sample size $n\in\{200,600,1000\}$. The exposures $\mathbf{x}$ were generated from a multivariate normal distribution with a moderate autoregressive correlation structure ($\rho$=0.5). Exposure-outcome associations were weak to moderate, with signal to noise ratios between 0.07 and 0.37, and we considered three scenarios that vary in their favorability to MIMs:
Scenario 1: $E[\bm{Y}\mid\bm{X}]=\left[\begin{array}[]{c}0.5f_{1}(\mathbf{x})\\ 0.5f_{1}(\mathbf{x})\\ 0.5f_{2}(\mathbf{x})\\ 0.3f_{2}(\mathbf{x})\end{array}\right]$ $f_{1}(\mathbf{x})=\mathbf{x}^{T}\bm{\alpha}_{1},$
$f_{2}(\mathbf{x})=\mathbf{x}^{T}\bm{\alpha}_{1}+(\mathbf{x}^{T}\bm{\alpha}_{2})^{2},$
@@
Scenario 2: $E[\bm{Y}\mid\bm{X}]=\left[\begin{array}[]{c}0.25f_{1}(\mathbf{x})\\ 0.25f_{1}(\mathbf{x})\\ 0.3f_{2}(\mathbf{x})\\ 0.3f_{2}(\mathbf{x})\end{array}\right]$ $f_{1}(\mathbf{x})=e^{0.5x_{1}}+\dfrac{1.5e^{1.2x_{2}}}{1+e^{1.2x_{2}}}-0.5x_{3}^{2},$
$f_{2}(\mathbf{x})=x_{9}-0.75x_{10}^{2}.$
Scenario 3: $E[\bm{Y}\mid\bm{X}]=\left[\begin{array}[]{c}f_{1}(\mathbf{x})\\ f_{1}(\mathbf{x})\\ \sqrt{0.5}f_{1}(\mathbf{x})\\ \sqrt{0.5}f_{1}(\mathbf{x})\end{array}\right]$ $f_{1}(\mathbf{x})=(\mathbf{x}^{T}\bm{\alpha}_{1})\times(\mathbf{x}^{T}\bm{\alpha}_{2}),$
@@
To each dataset we fit: (1) the proposed MIM separately for each outcome (MIM separate); (2) the proposed multivariate MIM, clustering across outcomes (MIM clustering); (3) Bayesian kernel machine regression (BKMR, bobb2015bayesian), a widely used Gaussian process approach for modelling non-linear and non-additive relationships with variable selection on the exposures; and (4) weighted quantile sum regression (WQS, carrico2015characterization), a popular mixtures approach that fits a linear index model on exposure quantiles.
@@
We also computed exposure importance metrics $\Phi_{kj}$ proposed in Section 4.1 for the MIM. To estimate the distribution of $x_{p}$ given $\mathbf{x}_{-p}$ we assumed normality and estimated the mean of $x_{p}$ given $\mathbf{x}_{-p}$ via random forests to allow for flexible associations between exposures. We report MSE for the posterior mean of the exposure-response surface averaged over 200 randomly chosen locations, and we also report the average exposure importance metrics.
@@
We report results for MSE in Fig. 3. In Scenario 1, the true exposure-response relationships fit into the index framework, and the MIM approaches naturally performed best. By contrast, Scenario 2 is favorable to BKMR, as the regression surfaces are sparse (70-80% of effects are null), and each non-null effect has a different functional form. As such, BKMR performed best here in small sample sizes. Surprisingly, however, the proposed MIM approaches began to outperform BKMR in larger samples, as BMKR tended to incorrectly exclude important exposures in some datasets. Scenario 3, characterized by many pairwise interactions, is not particularly favorable to either approach. Here the MIMs—despite having an additive structure—greatly outperformed the other approaches. Across all scenarios and sample sizes, MIM clustering successfully exploited shared structure across outcomes, and outperformed MIM separate (which did not borrow information across outcomes). Overall, the proposed approach fared well in a range of settings: under sparsity, it performed reasonably well, and when many exposures had small effects, it greatly outperformed competitors.
@@
Figure 3: Mean squared error for estimating $f^{*}(\mathbf{x})$ at 200 randomly chosen locations for $\mathbf{x}$ across all estimators, sample sizes, and scenarios. MSE values are scaled so that MIM clustering has an MSE of 1 in all scenarios.
@@
In Fig. 4, we see that estimation of exposure importance metrics is difficult in small sample sizes. In particular when $n=200$ we observe substantial overestimation in the highly sparse Scenario 2, in part because when $\Phi_{kp}=0$ it can never be underestimated (as $\Phi_{kp}\geq 0$ by definition). Ultimately, as sample size increased, estimation improved significantly in all three scenarios. This shows that the proposed exposure metrics can be reasonably estimated in order to provide additional interpretability to the proposed MIM framework.
@@ 6 NMMAPS Data Analysis
The National Morbidity, Mortality and Air Pollution Study (NMMAPS) collected time series data on daily mortality, air pollution and meteorological variables in Chicago between 1987 and 2000. We observe daily cardiovascular deaths, respiratory deaths and other non-accidental deaths, as well as median PM10 and O3 measurements. Full study details can be found in samet2000national; samet2000nationalb. Mortality outcomes were log transformed and standardized. Exposures were standardized, and we excluded one two week period in 1988 with a PM10 measurement 15 standard deviations above the mean, leaving a dataset of $n=5,100$ days, and we considered a maximum lag of $L=14$ days. We note that for this time series analysis, the unit of study is the day; for the $i^{th}$ day ($i=1,\dots,5100$), we observe outcomes $\mathbf{y}_{i}=(y_{i1},y_{i2},y_{i3})^{T}$, exposures $\mathbf{x}_{i1}=(x_{i1,1},\dots,x_{i1,14})^{T}$ and $\mathbf{x}_{i2}=(x_{i2,1},\dots,x_{i2,14})^{T}$, as well as covariates $\mathbf{z}_{i}$, including date and meteorological covariates. A preliminary additive model found no associations with same-day exposures, but this ignored delayed exposure effects; a lagged WQS (gennings2020lagged) analysis suggested negative associations with the exposure mixture (Appendix figures H19 & H20), though this ignored uncertainty in time-varying mixture weights.
@@
To quantify non-linear relationships between exposures and outcomes, we fit the proposed MV-DLNM with $K=3$ and $P=2$. Following chen2019distributed, we adjusted for day of week, seasonality, year, and mean daily temperature and dew point temperature per day and averaged over the prior three days, and we allowed different covariate associations for each outcome (see Appendix H). We compared the proposed MV-DLNM with cluster-inducing priors to separate DLNMs for each outcome, reflecting a standard approach that did not borrow strength across outcomes. We also considered the MV-DLNM without clustering as a middle ground between the two (with results in Appendix H).
The separate modelling approach indicated a strong non-linear association between O3 and cardiovascular deaths (Fig. 5b, left column). This relationship was mostly negative and linear for exposures less than 1SD above the mean but turned sharply positive beyond that. The relationship between PM10 was also non-linear, with associations first increasing, then decreasing, then increasing again for exposures above the mean. Associations with respiratory and other deaths were more uncertain, and intervals were wide enough to include strong positive or negative associations (Fig. 5, rows 2–3).
(a) PM10. Left: separate models; right: proposed MV-DLNM.
(b) O3. Left: separate models; right: proposed MV-DLNM.
@@
Figure 5: Mean difference in outcomes increasing standardized PM10 (panel a) and O3 (panel b) from their means. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths.
The proposed MV-DLNM also captured the non-linear O3–cardiovascular relationship (Fig. 5b, right), this time with statistically significantly more deaths at the 95th percentile vs. mean exposure. In contrast with the separate modelling strategy, other associations were close to null with much tighter intervals. These differences are explained by the posterior clustering probabilities (Appendix Fig. H21). Pairwise clustering probabilities for $\bm{\beta}_{kp}$ were all above 0.90. Those for $\bm{\theta}_{kp}$ were more modest, but five of six exposure-outcome pairs exhibited substantial clustering, with the exception of the O3 – cardiovascular relationship. The upshot is that the proposed approach exploited similarity in the weaker associations to gain efficiency while allowing the distinct non-linear relationship between O3 and cardiovascular deaths to shine through. We note that the PM10–cardiovascular relationship was flattened in this approach, possibly due to instability in the separate modelling estimates.
Overall mixture effects are often of interest, and in Fig. 6 we plot estimated mean differences when exposures are simultaneously set to different quantiles. Counterintuitively, the separate model for cardiovascular deaths showed a strong negative relationship with the overall mixture. The proposed MV-DLNM reflected a much more modest negative relationship, and exhibited a sharp upturn at the highest exposure levels. Separate model fits for the other outcomes exhibited wide intervals that could be consistent with null, positive, or negative relationships. The proposed MV-DLNM showed much tighter intervals, indicating more evidence of no relationship with the overall mixture.
Figure 6: Overall exposure effects: estimated mean outcomes setting all exposures to different quantiles (vs median). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is separate models; right column is proposed MV-DLNM.
To quantify temporal profiles we estimate mean outcome differences corresponding to an increase in a pollutant from 1SD below its mean to 1SD above at each lag $l$, holding exposures at all other lags constant (as in chen2019distributed). Estimates were similar across models here (Appendix figures H22H23), with weak associations across all three outcomes, in part because they represent single day changes in exposure while holding constant the other 13 days’ exposures (which are highly correlated). Nevertheless, separate models yielded intervals on average more than twice the width of those of the proposed MV-DLNM.
Ultimately, the proposed MV-DLNM fit better than modelling outcomes separately (WAIC 40159.8 vs 40282.6). We also fit the MV-DLNM without clustering, which had slightly worse fit (40161.4) and gave estimates more similar to separate model fits (Appendix H). The cluster-inducing priors enabled the MV-DLNM to capture the non-linear O3-cardiovascular relationship while exploiting similarity in the other weak associations to improve efficiency.
@@ 7 Discussion
Model (2) is a special case of our framework, but more flexible parameterizations could be considered, including non-parametric functional additive models (james2005functional; mclean2014functional), though the low signal of mixtures analyses motivates more structured approaches. A limitation of our MV-DLNM is that it assumes additivity. In Appendix J we discuss how our methods could be extended to non-additive settings; in our experience with environmental mixtures in settings with repeated exposures over time, however, we have found there is seldom enough signal to support identifying interactions, particularly in non-linear models. Nevertheless, alternative DLNMs could be considered, like the non-separable parameterizations of gasparrini2017penalized. In Appendix I we extend our approach to models of that type and demonstrate similar performance in simulations. Recently, li2025dynamic proposed a model that combines exposures into a single index and allows for a time-varying function of this index—though it restricts contributions of each exposure to be constant over time, and for each to share the same exposure-response function. Ultimately, our goal is not to argue the supremacy of any one specification but to present a flexible framework for borrowing strength across outcomes and exposures facilitated by our MV-DLNM parameterization. Ongoing work aims to develop methods for comparing and combining inferences from these different types of DLNMs.
In this paper we proposed a multivariate index model framework for modeling environmental mixtures with a novel co-clustering strategy to exploit similarity across exposures and outcomes, and we developed new exposure-importance metrics to aid interpretation. Though we highlighted the MV-DLNM for time-varying exposures, the proposed framework extends to other exposure structures. Rather than measuring $P$ pollutants at multiple times, for example, bauer2020associations analyzed five biomarkers (in hair, blood, urine, nails, and saliva) of four metals, for each of three outcomes via separate models. Our approach can naturally be applied to this setting to borrow strength in an analogous way to the MV-DLNM (omitting smoothness inducing priors over time). With univariate exposure measurements, our framework extends (multivariate) additive models by encouraging similar exposure-response curves across different exposure-outcome pairs (see Appendix A). Even if the outcome is univariate ($K=1$), the proposed approach may still offer a benefit by borrowing strength across a large number of exposures.
Data Availability Statement
The data analyzed in Section 6 are available in the dlnm package (dlnmpackage).
Supplementary Materials
R code for running the proposed methods are available at github.com/glenmcgee/MVmixture.